Consider the following discrete probability distribution:

X -0.99 0.48 0.71 1.4
P(X) 0.1 0.4 0.3 0.2

a) What is E[X]?

Round your answer to at least 3 decimal places.

b) What is Var[X]?

Round your answer to at least 3 decimal places.
Answer #1

a)
E[X] = sum(X*P(X))
= -0.99*0.1 + 0.48*0.4 + 0.71*0.3 + 1.4*0.2
= 0.586

b)
Var[X] = sum((X - E[X])^2*P(X))
= (-0.99 - 0.586)^2*0.1 + (0.48 - 0.586)^2*0.4 + (0.71 - 0.586)^2*0.3 + (-0.99 - 1.4)^2*0.2

= 1.400